OpenStudy (anonymous):

PLEEASSSEE HELP FOR MEDAL + FAN!!!! What is the slope? y = 5x - 4 The slope is ___.

mathslover (mathslover):

Compare this with general equation : y = mx + c

mathslover (mathslover):

where m = slope of line ; c = y - intercept ( a no.)

OpenStudy (anonymous):

graph it first and then count the spaces up and across and make it a fraction for the slope

OpenStudy (andrewhaze):

I thought it was y=mx+b

OpenStudy (anonymous):

okay what's going on??

OpenStudy (andrewhaze):

idk

OpenStudy (ghostgate):

Yeah, it's y = mx + b not y = mx + c.

OpenStudy (ghostgate):

The slope is 5. m = 5, b = y-intercept.

OpenStudy (anonymous):

so the slope is y=mx+b

OpenStudy (andrewhaze):

no

OpenStudy (anonymous):

oh wait never mind

OpenStudy (anwaarullah):

5 is slop and -4 is intercept

mathslover (mathslover):

Well, in India and some other countries we use, y = mx + c.... While in other countries, y = mx + b ...! it is just the notation that differs. Meaning remains the same.

mathslover (mathslover):

http://www.mathsisfun.com/equation_of_line.html#Countrynote ^ This is very interesting. I suggest you all to have a look at it. See how the symbols used differ from country to country. @anwaarullah @andrewhaze

OpenStudy (andrewhaze):

what math solver

OpenStudy (anonymous):

so who do i fan and give the medal to??

OpenStudy (andrewhaze):

I say @Ghostgate